# Learning Prompt with Distribution-Based Feature Replay for Few-Shot Class-Incremental Learning

### v1

@@ 4.1 Experimental Setup
Joint-LP 76.22 94.81 75.42
Figure 4: Accuracy curves of our LP-DiF and comparison with existing SOTA FSCIL methods on (a) CIFAR100 and (b) CUB200.
@@
Table 3: Ablation studies of our LP-DiF on mini-ImageNet. LP and OCD denote learning prompts and old-class distribution, respectively. RF and SF denote the real features of training images and synthesized features generated by VAE, respectively.
@@
Table 4: Comparison with other replay approaches on CUB-200 in terms of Avg. The results of “Randomly selection” are reports over $5$ runs with mean and standard deviations. ${}^{\dagger}$ means applying the replay technique proposed in iCaRL to CLIP + LP.
Methods $\bm{N}_{e}$ Storage Space Avg
@@
LP-DiF (Ours) - 0.78 MB 74.00
Figure 5: Accuracy curves of our LP-DiF and comparison with counterparts on (a) SUN-397 and (b) CUB200*.
4.2 Main Results
Comparison with State-of-The-Arts. We summarize the results of competing methods on mini-ImageNet in Table 1. Clearly, employing CLIP (baseline) [38, 29] for zero-shot evaluation alone outperforms all existing FSCIL methods by a large margin in terms of accuracy in each session and Average Accuracy (Avg). Naturally, it achieves a notably lower Performance Drop rate (PD). Our LP-DiF further achieves 16.19% ($77.57$% $\rightarrow$ $93.76$%) gains than the CLIP in terms of Avg, and shows comparable PD performance, i.e., $4.66$% vs. $4.56$%. As for the existing SOTA methods, e.g., NC-FSCIL [51], which presents the best Avg among all the SOTA methods, LP-DiF gains 25.94% improvements, i.e., 67.82% $\rightarrow$ 93.76%. Comparing with C-FSCIL [18], which presents the best PD. among the competing methods, LP-DiF gains 10.33% improvements, i.e., 14.99% $\rightarrow$ 4.66%. Fig. 4 shows the performance curves on CIFAR-100 and CUB-200. One can observe that LP-DiF consistently outperforms existing SOTA methods in all sessions for all benchmarks. These results clearly illustrate the superiority of our LP-DiF.
Comparison with Upper Bound. Assuming that the training set from each previous session is available, we can jointly train the prompts using these sets, thereby avoiding the issue of forgetting old information. In class-incremental learning, the above-mentioned setting can be considered as an upper bound, and serve as a reference for evaluating FSCIL method. Thus, we compare our LP-DiF with the upper bound (i.e. Joint-LP). As shown in Table 2, across the three benchmarks, the performance of our method is very close to the upper bound in terms of Avg, with the largest gap being only $1.42$% on CUB-200. The results indicate that our LP-DiF is highly effective in preventing catastrophic forgetting.
Figure 6: Ablation studies of our LP-DiF. (a) Comparison with the method of incorporating a Linear Classifier (LC) into a pre-trained image encoder for training. (b) Analysis of $M$. (c) Analysis of $B$ and $\lambda_{\text{o}}$ in terms of Avg on CUB-200.
More Challenging Benchmarks. On the three widely used benchmarks, the performance of our LP-DiF closely approaches the upper bound. To further assess our LP-DiF, we provide two more challenging benchmarks: SUN-397 and CUB-200* (refer to suppl. for details). For each challenging benchmark, we compare our LP-DiF with three distinct approaches: zero-shot evaluation using CLIP (baseline), Joint-LP (upper bound), and BiDistFSCIL [56] (SOTA open-source method). The corresponding performance curves are depicted in Fig. 5. Overall, on both SUN-397 and CUB-200*, our LP-DiF method 1) significantly surpasses both CLIP and BiDistFSCIL, and 2) attains performance levels that are very close to the respective upper bounds. The results show that our LP-DiF remains very effective on these challenging situations, including those with a larger number of classes and extended session lengths, e.g., $397$ classes across $21$ sessions in SUN-397, as well as in those without a base session, exemplified by CUB-200*.
4.3 Ablation Study
Analysis of Key Components. Our proposed method involves prompt tuning on CLIP to adapt the knowledge from each incremental session. It also constructs feature-level distributions to preserve old knowledge, thereby achieving resistance to catastrophic forgetting. To investigate the effect of the key components in our method, i.e., CLIP, prompt learning (LP), the distribution estimated by real features (RF) of training images, and the synthesized features (SF), we summarized the performance of each component on mini-ImageNet in Table 3. As illustrated, employing the LP technique noticeably improves performance across each session, ultimately resulting in a superior $11.74$% performance, i.e., from $77.57$% $\rightarrow$ $89.31$% in terms of average performance (refer to the second row of Table 3). However, solely implementing LP causes higher PD than CLIP, e.g., $4.56$% $\rightarrow$ $11.58$%, due to the forgetting of old knowledge during learning in new sessions. Additionally, as mentioned in Sec.3.2, the old-class distribution (OCD) is effective in tackling the forgetting problem. Note that the distribution is estimated by real features (RF) and the synthesized features (SF) generated by VAE. So we conducted separate evaluations to assess the effect of these two types of “features”. Concretely, using only RF to estimate the old-class distribution can improve the Avg by $4.11$%, i.e., $89.31$% $\rightarrow$ 93.42%, and reduce the PD. by 6.12%, i.e., $11.58$% $\rightarrow$ 5.46%, (see the third row of Table 3). Using only SF can also improve Avg and reduce PD, however, its effectiveness is marginally inferior to using only RF (refer to the fourth row of Table 3). Finally, using both types of “features” can further improve the performance, which surpasses the outcomes achieved by using either RF or SF alone (see the last row of Table 3). Thus, although LP can enable the model to effectively capture the knowledge of each session and improve performance, it still leads to catastrophic forgetting, while OCD can effectively prevent this issue.
Analysis of $M$. As mentioned in Sec. 3.2, $M$ is the number of synthetic features which participate in estimating $\bm{\mu}_{c}$ and $\bm{\sigma}_{c}$ of old-class distribution. The value of $M$ will affect the accuracy of the estimated distribution, thus influencing the performance of the model. Therefore, we test the effect of $M$ on the performance. Fig. 6 (b) shows the results on the three widely used benchmarks. Clearly, when $M$ increases from 00, the Avg gradually improves. Nonetheless, when $M$ continues to increase, the Avg decreases slightly, possibly ascribing to that too many synthesized features can cause the estimated distribution to overly skew towards the distribution of the synthesized features. In summary, $M=10$ (on mini-ImageNet) or $M=15$ (on CIFAR-$100$ and CUB-$200$) are the best choices for performance.
Old-Class Distribution vs. Image Exemplar. To further validate its efficacy in avoiding catastrophic forgetting, we compare our method with other replay-based approaches tailored for learning prompts, i.e., 1) randomly selecting $N_{e}$ images of per old class as exemplars; 2) adopting the replay strategy in iCaRL [30], specifically choosing $N_{e}$ images for each old class based on the proximity to the mean feature. In addition, we execute the random selection approach five times, each with a different random seed, to reduce the uncertainty. The average results with necessary storage space for replay on CUB-200 are shown in Table 4, where CLIP + LP indicates the baseline without replay of old classes. Obviously, our method exhibits the best performance and lowest storage space in comparison to the two counterparts under various $N_{e}$. Especially, compared with ${}^{\dagger}$ under $N_{e}=4$, LP-DiF shows a comparable performance while only requiring about $\textbf{0.01}\%$ storage space. This underscores that our pseudo-feature replay technique can effectively combat catastrophic forgetting under conditions of light storage overhead.
5 Conclusion
In this paper, we studied the FSCIL problem by introducing V-L pretrained model, and proposed Learning Prompt with Distribution-based Feature replay (LP-DiF). Specifically, prompt tuning is involved to adaptively capture the knowledge of each session. To alleviate catastrophic forgetting, we established a feature-level distribution for each class, which is estimated by both real features of training images and synthesized features generated by a VAE decoder. Then, pseudo features are sampled from old-class distributions, and combined with the training set of current session to train the prompts jointly. Extensive experiments demonstrate that our LP-DiF achieves the new state-of-the-art in the FSCIL task.
Appendix
Contents
The following items are included in the appendix:
  • Details of Selected Benchmarks in Sec. 6.
  • Quantitative Comparison with SOTAs on CUB-200 and CIFAR-100 in Sec. 7.
  • Comparison with Pre-trained Model-based Standard CIL Methods in Sec. 8.
  • Decomposing the Performance of the Base Class and the Incremental Class in Sec. 9.
  • Analysis of Various Shot Numbers in Sec. 10.
6 Details of Benchmarks
Following the mainstream benchmark settings [56], we conduct our experiments on three datasets, i.e., CIFAR-100 [25], mini-ImageNet [40] and CUB-200 [41], to evaluate our LP-DiF.
Additionally, this paper also proposes two more challenging benchmarks for FSCIL, i.e., SUN-397 [48] and $*{}^{*}$:
  • $*{}^{*}$ is a variant of CUB-200 but excludes the base session. We evenly divide the total $200$ classes into $20$ incremental sessions, with each session containing $10$ categories. Following the previous study [56], there are $100$ classes in the base session, and the remaining classes will be divided into $10$ incremental sessions, with each incremental session comprising $10$ classes. We use it to evaluate whether our method works in scenarios without the base session.
For the above five benchmarks, the classes in the base session have a large amount of training data, while each class in every incremental session has only $5$ training images. Tab. 5 summarizes the details of each selected benchmark.
Table 5: Details of selected benchmarks. The first three lines are commonly used benchmarks, while the last two lines are the more challenging benchmarks proposed in this paper. $\bm{|\mathcal{C}^{\text{All}}|}$, $\bm{|\mathcal{C}^{\text{Base}}|}$ and $\bm{|\mathcal{C}^{\text{Inc}}|}$ denotes the total number of classes, the number of classes in base session, and the number of classes in each incremental session respectively. #Base and #Inc denote the number of base sessions and the incremental session respectively. Shot denotes the number of training images of each incremental session. $*{}^{*}$ represents a variant version.
CIFAR-100 [25] 100 60 5 1 8 5
mini-ImageNet [32] 100 60 5 1 8 5
CUB-200 [41] 200 100 10 1 10 5
SUN-397 [48] 397 197 10 1 20 5
CUB-200* [41] 200 0 10 0 20 5
Table 6: Comparison with state-of-the-art FSCIL methods on CUB-200, where the rows filled with gray indicate the existing SOTA FSCIL methods. “Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance.
Methods Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
0 1 2 3 4 5  6 7 8 9 10
TOPIC [37] $68.68$ $62.49$ $54.81$ $49.99$ $45.25$ $41.40$ $38.35$ $35.36$ $32.22$ $28.31$ $26.26$ $43.92$ $42.42$
CEC [54] $75.85$ $71.94$ $68.50$ $63.50$ $62.43$ $58.27$ $57.73$ $55.81$ $54.83$ $53.52$ $52.28$ $61.33$ $23.57$
Replay [27] $75.90$ $72.14$ $68.64$ $63.76$ $62.58$ $59.11$ $57.82$ $55.89$ $54.92$ $53.58$ $52.39$ $61.52$ $23.51$
MetaFSCIL [9] $75.90$ $72.41$ $68.78$ $64.78$ $62.96$ $59.99$ $58.30$ $56.85$ $54.78$ $53.82$ $52.64$ $61.93$ $23.26$
FACT [57] $75.90$ $73.23$ $70.84$ $66.13$ $65.56$ $62.15$ $61.74$ $59.83$ $58.41$ $57.89$ $56.94$ $64.42$ $18.96$
FCIL [15] $78.70$ $75.12$ $70.10$ $66.26$ $66.51$ $64.01$ $62.69$ $61.00$ $60.36$ $59.45$ $58.48$ $65.70$ $20.22$
GKEAL [62] $78.88$ $75.62$ $72.32$ $68.62$ $67.23$ $64.26$ $62.98$ $61.89$ $60.20$ $59.21$ $58.67$ $66.35$ $20.21$
NC-FSCIL [51] $80.45$ $75.98$ $72.30$ $70.28$ $68.17$ $65.16$ $64.43$ $63.25$ $60.66$ $60.01$ $59.44$ $67.28$ $21.01$
BiDistFSCIL [56] $79.12$ $75.37$ $72.80$ $69.05$ $67.53$ $65.12$ $64.00$ $63.51$ $61.87$ $61.47$ $60.93$ $67.34$ $18.19$
SAVC [36] $81.85$ $77.92$ $74.95$ $70.21$ $69.96$ $67.02$ $66.16$ $65.30$ $63.84$ $63.15$ $62.50$ $69.35$ $19.35$
F2M [33] $81.07$ $78.16$ $75.57$ $72.89$ $70.86$ $68.17$ $67.01$ $65.26$ $63.36$ $61.76$ $60.26$ $69.49$ $20.81$
CLIP (Baseline[29] $65.54$ $62.91$ $61.54$ $57.75$ $57.88$ $57.89$ $56.62$ $55.40$ $54.20$ $54.23$ $55.06$ $58.09$ 10.48
LP-DiF (Ours) 83.94 80.59 79.17 74.30 73.89 73.44 71.60 70.81 69.08 68.74 68.53 74.00 $15.41$
[…]
Table 7: Comparison with state-of-the-art FSCIL methods on CIFAR-100, where the rows filled with gray indicate the existing SOTA FSCIL methods. “Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance.
Methods Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
0 1 2 3 4 5  6 7 8
TOPIC [37] $64.10$ $55.88$ $47.07$ $45.16$ $40.11$ $36.38$ $33.96$ $31.55$ $29.37$ $42.62$ $34.73$
F2M [33] $64.71$ $62.05$ $59.01$ $55.58$ $52.55$ $49.96$ $48.08$ $46.28$ $44.67$ $53.65$ $20.04$
CEC [54] $73.07$ $68.88$ $65.26$ $61.19$ $58.09$ $55.57$ $53.22$ $51.34$ $49.14$ $59.53$ $23.93$
Replay [27] $74.40$ $70.20$ $66.54$ $62.51$ $59.71$ $56.58$ $54.52$ $52.39$ $50.14$ $60.78$ $24.26$
MetaFSCIL [9] $74.50$ $70.10$ $66.84$ $62.77$ $59.48$ $56.52$ $54.36$ $52.56$ $49.97$ $60.79$ $24.53$
GKEAL [62] $74.01$ $70.45$ $67.01$ $63.08$ $60.01$ $57.30$ $55.50$ $53.39$ $51.40$ $61.35$ $22.61$
C-FSCIL [18] $77.47$ $72.40$ $67.47$ $63.25$ $59.84$ $56.95$ $54.42$ $52.47$ $50.47$ $61.64$ $27.00$
FCIL [15] $77.12$ $72.42$ $68.31$ $64.47$ $61.18$ $58.17$ $56.06$ $54.19$ $52.02$ $62.66$ $25.10$
FACT [57] $78.22$ $72.40$ $68.57$ $64.73$ $61.40$ $58.57$ $56.30$ $53.83$ $51.72$ $62.86$ $26.50$
SAVC [36] $78.77$ $73.31$ $69.31$ $64.93$ $61.70$ $59.25$ $57.13$ $55.19$ $53.12$ $63.63$ $25.65$
BiDistFSCIL [56] $79.45$ $75.38$ $71.84$ $67.95$ $64.96$ $61.95$ $60.16$ $57.67$ $55.88$ $66.14$ $23.57$
NC-FSCIL [51] $82.52$ $76.82$ $73.34$ $69.68$ $66.19$ $62.85$ $60.96$ $59.02$ $56.11$ $67.50$ $26.41$
CLIP (Baseline[29] $74.44$ $72.96$ $72.21$ $70.49$ $70.18$ $70.00$ $69.81$ $69.23$ $68.37$ $70.86$ 6.07
LP-DiF (Ours) 80.23 77.75 76.78 74.62 74.03 73.87 73.84 72.96 72.02 75.12 $8.21$
[…]
7 Quantitative comparison with SOTAs on CUB-200 and CIFAR-100
We have shown the accuracy comparison curves with the state-of-the-art methods on CIFAR-100 and CUB-200 in Fig. 4 of our main paper. Here, we present details comparison results with specific values on these two benchmarks, shown in Tab. 6 and Tab. 7. Overall, the performance of our LP-DiF can be summarized in three points. 1) There are significant improvements compared to CLIP (baseline) in terms of Avg (i.e., 15.91% and 4.26% improvements on CUB-200 and CIFAR-100 respectively). 2) Compared to existing SOTA methods, LP-DiF achieves a higher Avg and lower PD. 3) Both Avg and PD of LP-DiF are very close to that of Joint-LP (upper bound), i.e., lower only 1.88% Avg and 3.35% PD on CUB-200, 1.10% Avg and 0.81% PD on CIFAR-100. These results clearly illustrate the effectiveness of our LP-DiF.
8 Comparison with Pre-trained Models-based Standard CIL Methods
To further demonstrate the superiority of our method, we compare it with several recent standard CIL [59] methods and variant it to pre-trained model-based methods: L2P [46], DualPrompt [45] and AttriCLIP [43]. We reproduce these three approaches on CIFAR-100, CUB-200, and mini-ImageNet and evaluate them under FSCIL protocol respectively. As shown in Tab. 8, our LP-DiF outperforms these methods by a large margin in terms of Avg across all three benchmarks. We also find that these methods based on pre-trained models underperform BiDistFSCIL [56] on CIFAR-100 and CUB-200. This indicates that these methods are not advantageous for the FSCIL setting, further underscoring the effectiveness and significance of our method.
Table 8: Comparison with standard CIL methods based on pre-trained models on the three common benchmarks in terms of Avg. ${}^{\ddagger}$ indicates our reproduction.
Methods CIFAR-100 mini-ImageNet CUB-200
BiDistFSCIL [56] $66.14$ $62.04$ $67.34$
${}^{\ddagger}$ [46] $61.77$ $75.68$ $56.95$
${}^{\ddagger}$ [45] $63.50$ $76.61$ $62.32$
${}^{\ddagger}$ [43] $59.24$ $81.74$ $47.81$
LP-DiF (Ours) 75.12 93.76 74.00
Figure 7: Decomposing the performance of the base class and the incremental class. The performance is evaluated by the model from the last session on CUB-200.
9 Decomposing the Performance of the Base Class and the Incremental Class
Following previous studies [56, 54, 57], in this section, we decompose the accuracy, respectively analyzing the effectiveness of our LP-DiF for the classes in the base session (i.e., base class) and for the classes in incremental sessions (i.e., incremental class), to evaluate if our method performs well on both base and incremental classes. We report the comparison results in terms of individual accuracy of base and novel classes, as well as their harmonic mean, in the last session on CUB-200. Fig. 7 shows that our LP-DiF outperforms the second best method on base class (i.e., FACT) by $3.5$%, while outperforms the second best method on incremental class (i.e., BiDistFSCIL) by $9.9$%. Finally, the superior harmonic mean demonstrates our achievement of an enhanced balance between base and novel classes.
@@ 9 Decomposing the Performance of the Base Class and the Incremental Class
10 Analysis Various Shot Numbers

### v3

@@ IV-A Datasets and Metrics
TABLE VII: Comparison with SOTA FSCIL methods combined with CLIP on three common benchmarks in terms of Avg. We replace the original backbone of these methods with CLIP’s image encoder (ViT-B/16), and we use the text encoding generated by CLIP’s text encoder for each category as initialization for the classification layer parameters of these methods.
Methods CIFAR-100 mini-ImageNet CUB-200
SAVC [39] + CLIP $68.46$ $86.81$ $71.66$
BiDistFSCIL [60] + CLIP $69.40$ $84.63$ $70.95$
LP-DiF (Ours) 75.12 93.76 74.00
IV-C Main Results
Comparison with State-of-The-Arts. We summarize the results of competing methods on mini-ImageNet in Table I. Clearly, employing CLIP (baseline) [41, 32] for zero-shot evaluation alone outperforms all existing FSCIL methods by a large margin in terms of accuracy in each session and Average Accuracy (Avg). Naturally, it achieves a notably lower Performance Drop rate (PD). Our LP-DiF further achieves 16.19% ($77.57$% $\rightarrow$ $93.76$%) gains than the CLIP in terms of Avg, and shows comparable PD performance, i.e., $4.66$% vs. $4.56$%. As for the existing SOTA methods, e.g., NC-FSCIL [54], which presents the best Avg among all the SOTA methods, LP-DiF gains 25.94% improvements, i.e., 67.82% $\rightarrow$ 93.76%. Comparing with C-FSCIL [21], which presents the best PD. among the competing methods, LP-DiF gains 10.33% improvements, i.e., 14.99% $\rightarrow$ 4.66%. Tab. IV highlights the comparison with the replay-based FSCIL method. Note that existing replay-based methods directly store old-class images, while our method only requires storing one mean vector and one variance vector (diagonal elements of the covariance matrix) for each old class. Generally, LP-DiF significantly outperforms others in terms of performance while costing the least amount of storage space. Tab. III and Tab. V shows the comparison results on CUB-200 and CIFAR-100. Overall, the performance of our LP-DiF can be summarized in two points. 1) There are significant improvements compared to CLIP (baseline) in terms of Avg (i.e., 15.91% and 4.26% improvements on CUB-200 and CIFAR-100 respectively). 2) Compared to existing SOTA methods, LP-DiF achieves a higher Avg and lower PD. Moreover, considering that our method use CLIP as backbone, which is more stronger than those existing FSCIL methods whose backbone are ResNet, we replaced the backbone of these methods with CLIP to make a fairer comparison. Specifically, we replace the original backbone of these methods with CLIP’s image encoder (ViT-B/16), and we use the text encoding generated by CLIP’s text encoder for each category as initialization for the classification layer parameters of these methods. Tab. VII shows that the performance of the existing SOTA FSCIL methods combined with CLIP are still lower than our LP-DiF. These above results clearly illustrate the superiority of our LP-DiF.
Comparison with Upper Bound. Assuming that the training set from each previous session is available, we can jointly train the prompts using these sets, thereby avoiding the issue of forgetting old information. In class-incremental learning, the above-mentioned setting can be considered as an upper bound, and serve as a reference for evaluating FSCIL method. Thus, we compare our LP-DiF with its upper bound (i.e. Joint-LP). As shown in the last row of Tab. I, Tab. III and Tab. V across the three benchmarks, the performances of our method are very close to the upper bounds in terms of Avg, with the largest gap being only $1.05\%$, $1.42\%$ and $1.10\%$ on mini-ImageNet, CUB-200 and CIFAR-100, respectively. The results indicate that our LP-DiF is highly effective in preventing catastrophic forgetting. It is noted that NC-FSCIL achieves higher accuracy than both ours and Joint-LP. The architectures of NC-FSCIL, which is ResNet-based method, and Joint-LP, which is CLIP-based method, are different; NC-FSCIL trains all layers of the model during the base session, whereas Joint-LP only train the prompt. Therefore, it is acceptable that NC-FSCIL outperforms Joint-LP in session 0.
Comparison with Pre-trained Models-based Standard CIL Methods. To further demonstrate the superiority of our method, we compare it with several recent standard CIL [62] methods which also utilize pre-trained models: L2P [49], DualPrompt [48] and AttriCLIP [45]. The L2P and DualPrompt are based on a pretrained ViT and learn the visual prompts to solve CIL problems, while AttriCLIP builds on CLIP and training different text prompts to encode different knowledge. We reproduce these three approaches on CIFAR-100, CUB-200, and mini-ImageNet and evaluate them under FSCIL protocol respectively. As shown in Tab. VI, our LP-DiF outperforms these methods by a large margin in terms of Avg across all three benchmarks. We also find that these methods based on pre-trained models underperform BiDistFSCIL [60] on CIFAR-100 and CUB-200. This indicates that these methods are not advantageous for the FSCIL setting, further underscoring the effectiveness and significance of our method.
More Challenging Benchmarks. On the three widely used benchmarks, the performance of our LP-DiF closely approaches the upper bound. To further assess our LP-DiF, we provide two more challenging benchmarks: SUN-397 and CUB-200*. For each challenging benchmark, we compare our LP-DiF with three distinct approaches: zero-shot evaluation using CLIP (baseline), Joint-LP (upper bound), and BiDistFSCIL [60] (SOTA open-source method). The corresponding performance curves are depicted in Fig. 4. Overall, on both SUN-397 and CUB-200*, our LP-DiF method 1) significantly surpasses both CLIP and BiDistFSCIL, and 2) attains performance levels that are very close to the respective upper bounds. The results show that our LP-DiF remains very effective on these challenging situations, including those with a larger number of classes and extended session lengths, e.g., $397$ classes across $21$ sessions in SUN-397, as well as in those without a base session, exemplified by CUB-200*.
@@ IV-C Main Results
TABLE VIII: Ablation studies of our LP-DiF on mini-ImageNet. LP and OCD denote learning prompts and old-class distribution, respectively. RF and SF denote the real features of training images and synthesized features generated by VAE, respectively.
@@
TABLE IX: Ablation studies of our LP-DiF on CUB-200. LP and OCD denote learning prompts and old-class distribution, respectively. RF and SF denote the real features of training images and synthesized features generated by VAE, respectively.
RF SF 0 1 2 3 4 5 6 7 8 9 10
$65.54$ $62.91$ $61.54$ $57.75$ $57.88$ $57.89$ $56.62$ $55.40$ $54.20$ $54.23$ $55.06$ $58.09$ 10.48
83.94 $78.32$ $75.10$ $70.62$ $70.75$ $68.09$ $65.69$ $64.55$ $62.47$ $61.94$ $61.96$ $70.71$ $21.98$
83.94 80.59 $78.83$ $73.66$ $73.24$ $72.54$ $70.57$ $69.72$ $68.88$ $67.86$ $67.90$ $73.43$ $16.04$
83.94 80.59 $78.41$ $72.65$ $72.76$ $71.25$ $69.86$ $67.99$ $67.20$ $66.73$ $66.88$ $72.56$ $17.06$
83.94 80.59 79.17 74.30 73.89 73.44 71.60 70.81 69.08 68.74 68.53 74.00 $15.41$
TABLE X: Ablation studies of our LP-DiF on CIFAR-100. LP and OCD denote learning prompts and old-class distribution, respectively. RF and SF denote the real features of training images and synthesized features generated by VAE, respectively.
RF SF 0 1 2 3 4 5 6 7 8
$74.44$ $72.96$ $72.21$ $70.49$ $70.18$ $70.00$ $69.81$ $69.23$ $68.37$ $70.86$ 6.07
80.23 75.8175.81{75.81}75.81 75.0375.03{75.03}75.03 71.6571.65{71.65}71.65 71.6771.67{71.67}71.67 70.9470.94{70.94}70.94 70.4870.48{70.48}70.48 70.0170.01{70.01}70.01 69.5469.54{69.54}69.54 72.8172.81{72.81}72.81 $10.69$
80.23 77.75 76.84 74.4074.40{74.40}74.40 73.8173.81{73.81}73.81 73.2473.24{73.24}73.24 73.6973.69{73.69}73.69 72.5272.52{72.52}72.52 71.6071.60{71.60}71.60 74.8974.89{74.89}74.89 $8.63$
80.23 77.75 75.6375.63{75.63}75.63 73.7573.75{73.75}73.75 73.0973.09{73.09}73.09 72.3672.36{72.36}72.36 72.3172.31{72.31}72.31 71.8471.84{71.84}71.84 70.7670.76{70.76}70.76 74.1974.19{74.19}74.19 $9.47$
80.23 77.75 $76.78$ 74.62 74.03 73.87 73.84 72.96 72.02 75.12 $8.21$
Figure 5: Ablation studies of our LP-DiF. (a) Comparison with the method of incorporating a Linear Classifier (LC) into a pre-trained image encoder for training on three common benchmarks. (b) Analysis of $M$ on three common benchmarks. (c) Analysis of $B$ and $\lambda_{\text{o}}$ in terms of Avg on CUB-200.
Figure 6: Decomposing the performance of the base class and the incremental class. Their harmonic mean are also reported. The performance is evaluated by the model from the last session on CUB-200.
IV-D Ablation Studies and Analysis
Analysis of Key Components. Our proposed method involves prompt tuning on CLIP to adapt the knowledge from each incremental session. It also constructs feature-level distributions to preserve old knowledge, thereby achieving resistance to catastrophic forgetting. To investigate the effect of the key components in our method, i.e., CLIP, prompt learning (LP), the distribution estimated by real features (RF) of training images, and the synthesized features (SF), we summarized the performance of each component on three common FSCIL benchmark in Tab. VIII, Tab. IX and Tab. X. Take the results on mini-ImageNet as an example, as illustrated in Tab. VIII, employing the LP technique noticeably improves performance across each session, ultimately resulting in a superior $11.74$% performance, i.e., from $77.57$% $\rightarrow$ $89.31$% in terms of average performance (refer to the second row of Table VIII). However, solely implementing LP causes higher PD than CLIP, e.g., $4.56$% $\rightarrow$ $11.58$%, due to the forgetting of old knowledge during learning in new sessions. Additionally, as mentioned in Sec.III-B, the old-class distribution (OCD) is effective in tackling the forgetting problem. Note that the distribution of each incremental session is estimated by real features (RF) and the synthesized features (SF) generated by VAE. So we conducted separate evaluations to assess the effect of these two types of “features”. Concretely, using only RF to estimate the old-class distribution can improve the Avg by $4.11$%, i.e., $89.31$% $\rightarrow$ 93.42%, and reduce the PD. by 6.12%, i.e., $11.58$% $\rightarrow$ 5.46%, (see the third row of Table VIII). Using only SF for each incremental session can also improve Avg and reduce PD, however, its effectiveness is marginally inferior to using only RF (refer to the fourth row of Table VIII). Finally, using both types of “features” can further improve the performance, which surpasses the outcomes achieved by using either RF or SF alone (see the last row of Table VIII). Thus, although LP can enable the model to effectively capture the knowledge of each session and improve performance, it still leads to catastrophic forgetting, while OCD can effectively prevent this issue.
Analysis of $M$. As mentioned in Sec. III-B, $M$ is the number of synthetic features which participate in estimating $\bm{\mu}_{c}$ and $\bm{\sigma}_{c}$ of old-class distribution. The value of $M$ will affect the accuracy of the estimated distribution, thus influencing the performance of the model. Therefore, we test the effect of $M$ on the performance. Fig. 5 (b) shows the results on the three widely used benchmarks. Clearly, when $M$ increases from 00, the Avg gradually improves. Nonetheless, when $M$ continues to increase, the Avg decreases slightly, possibly ascribing to that too many synthesized features can cause the estimated distribution to overly skew towards the distribution of the synthesized features. In summary, $M=10$ (on mini-ImageNet) or $M=15$ (on CIFAR-$100$ and CUB-$200$) are the best choices for performance.
Figure 7: Analysis the effectiveness of VAE on mini-ImageNet. We calculate the difference between the estimated GD of each incremental class and corresponding reference GD (ideal GD, i.e., computed using the full training images of that class) by KL divergence. The red dots represents KL divergence between ideal GD and GD estimated by using both SF and RF. The blue dots represents KL divergence between ideal GD and GD estimated by using only RF.
@@ IV-D Ablation Studies and Analysis
TABLE XI: Comparison with other replay approaches on CUB-200 in terms of Avg. The results of “Randomly selection” are reports over $5$ runs with mean and standard deviations. iCaRL means applying the replay technique proposed in iCaRL to CLIP + LP.
Methods $\bm{N}_{e}$ Disk Space Avg
@@
LP-DiF (Ours) - 0.22 MB 74.00
@@
Old-Class Distribution vs. Image Exemplar. To further validate its efficacy in avoiding catastrophic forgetting, we compare our method with other replay-based approaches tailored for learning prompts, i.e., 1) randomly selecting $N_{e}$ images of per old class as exemplars; 2) adopting the replay strategy in iCaRL [33], specifically choosing $N_{e}$ images for each old class based on the proximity to the mean feature. In addition, we execute the random selection approach five times, each with a different random seed, to reduce the uncertainty. The average results with necessary storage space for replay on CUB-200 are shown in Table XI, where CLIP + LP indicates learning prompts sequentially across each incremental session without replay of old classes. (e.g. the second row in Tab. IX). Obviously, our method exhibits the best performance and lowest storage space in comparison to the two counterparts under various $N_{e}$. Especially, compared with iCaRL under $N_{e}=4$, LP-DiF shows a comparable performance while only requiring about $\textbf{0.002}\%$ storage space (thanks to the fact that we only store two vectors for each old class). This underscores that our pseudo-feature replay technique can effectively combat catastrophic forgetting under conditions of light storage overhead.
Decomposing the Performance of Base and Incremental Classes. Following previous studies [60, 56, 61], in this section, we decompose the accuracy, respectively analyzing the effectiveness of our LP-DiF for the classes in the base session (i.e., base class) and for the classes in incremental sessions (i.e., incremental class), to evaluate if our method performs well on both base and incremental classes. We report the comparison results in terms of individual accuracy of base and novel classes, as well as their harmonic mean, in the last session on CUB-200. Fig. 6 shows that our LP-DiF outperforms the second best method on base class (i.e., FACT) by $3.5$%, while outperforms the second best method on incremental class (i.e., BiDistFSCIL) by $9.9$%. Finally, the superior harmonic mean demonstrates our achievement of an enhanced balance between base and novel classes.